# Topology-guided Hypergraph Transformer Network: Unveiling Structural Insights for Improved Representation

### v2

@@ 4.2. Result
Hypergraphs, known for their adeptness at representing intricate higher-order relationships, have become invaluable in numerous scientific endeavors. The proposed THTN model builds on this, representing communities as hyperedges and thus forming a hypergraph that captures intricate node relations. In contrast to existing models that do not consider local connection information during hypergraph construction, THTN leverages GCN to preserve this vital information. Additionally, THTN integrates unique encoding modules to learn structural-spatial information and a structure-aware self-attention module. These modules enable THTN to produce robust node representations, recognizing key nodes and hyperedges from both structural and semantic perspectives.
4.2.1. Case study - Examining the impact of local and global structural knowledge
@@
To delve deeper into the roles of local and global structural knowledge, we train THTN excluding local structural components (specifically, $lse$, $ce$, $lc$, and $kc$) and omitting global structural elements (namely, $ue$, $hd$, and $cc$) on all dataset. Outcomes in Table 3 illustrate that local structural information significantly influences THTN efficacy more than its global counterpart. Given that hypergraphs inherently harness higher-order global structural details from the graph, the omission of global encodings in THTN does not critically hinder its performance. On the other hand, converting a graph to a hypergraph can result in a loss of local structural details, underscoring the importance of integrating local structural insights into hypergraph learning.
4.2.2. Case study - Examining the impact of different design schemes
@@
To better understand the impact of each design scheme, we conduct ablation studies on a small and a large dataset, Cora and PubMed, respectively. We consider running THTN without (W/O) considering a specific design scheme each time. The results are presented in Table 4, which implies the impact of the individual design scheme on the model performance. For instance, removing positional encoding ($pe$) results in an accuracy of 85.51% for the Cora dataset and 86.22% for the PubMed dataset. This table shows that for both datasets, the greatest performance degradation occurs when the local structure encoding ($lse$) is excluded from the input node feature. When we create hyperedge from a community, we lose the local connection information between the nodes, which might be very important for the hyperedges. $lse$ preserves local connection information by applying GCN to the input graph. The significant decrease in performance without $lse$ highlights its crucial contribution to the model’s overall performance.
4.2.3. Case Study - Examining the Impact of Feature Normalization and Concatenation
@@
As illustrated in Equation 4, our method combines various structural and spatial features (i.e., encodings) with the initial features of the nodes by summing them. To investigate the effect of normalizing these features before combining them with the node’s initial features, we conduct an additional experiment. In this experiment, we normalize the features prior to their integration. The results of this experiment are presented in Table 5 and are labeled as THTNN. Additionally, we perform another experiment to evaluate the impact of concatenating the features instead of summing them. The outcomes of this approach are also shown in Table 5 and are referred to as THTNC. A comparative analysis of the different versions of THTN presented in the table reveals that each version performs better on different datasets. However, the original THTN method generally outperforms both the normalized variant (THTNN) and the concatenated variant (THTNC).
@@
Model Cora Citeseer PubMed DBLP LastFMAsia
THTN 88.48 78.09 89.01 84.38 88.58
THTNN 86.88 76.23 88.41 85.46 87.26
THTNC 89.13 77.83 87.18 79.32 87.78
4.2.4. Case study - Examining the impact of global nodes
We investigate the effect of global nodes ($n_{g}$) on the model’s performance, illustrated in Figure 3. As per Figure 3 (a), the Cora dataset reveals a rising trend in accuracy as the global nodes $n_{g}$ increase, peaking at 88.48% with $n_{g}=3$, followed by a decline. Analogously, Figure 3 (b) presents the Citeseer dataset, where an increment in global nodes prompts an upsurge in accuracy, achieving a maximum of 78.09% at $n_{g}=1$ before the accuracy trend reverses. Similarly, we get the best performance for the PubMed, DBLP, and LastFMAsia datasets for $n_{g}=4$, $n_{g}=5$, and $n_{g}=4$, respectively. Overall, this trend suggests that an optimal number of global nodes can effectively incorporate relevant global information. However, exceeding this optimal number may introduce excessive parameters and reduced generalization ability. Thus, selecting an appropriate number of global nodes is crucial for optimal performance.
Figure 3. The performance of THTN with different numbers of global nodes ($n_{g}$).
@@ 5. Conclusion
In this paper, we introduce the Topology-guided Hypergraph Transformer Network (THTN) as a novel hypergraph neural network model for node classification. While existing hypergraph transformer models mainly rely on semantic feature-based self-attention, they ignore topological information. The THTN has addressed this limitation by presenting a structural and spatial encoding module along with a structure-aware self-attention mechanism. By considering both semantic and structural information, THTN provides a comprehensive approach to node representation learning. Experimental findings underscore THTN’s superior node classification capabilities relative to benchmarks, highlighting its potential to revolutionize hypergraph analytics.

### v4

@@ 4.2. Comparison with Baseline Methods
Hypergraphs, known for their adeptness at representing intricate higher-order relationships, have become invaluable in numerous scientific endeavors. The proposed THTN model builds on this, representing communities as hyperedges and thus forming a hypergraph that captures intricate node relations. In contrast to existing models that do not consider local connection information during hypergraph construction, THTN leverages GCN to preserve this vital information. Additionally, THTN integrates unique encoding modules to learn structural-spatial information and a structure-aware attention module. These modules enable THTN to produce robust node representations, recognizing key nodes and hyperedges from both structural and semantic perspectives.
4.3. Case Studies
In this section, we present several case studies.
@@ 4.3. Case Studies
i. Impact of local/global structural knowledge To delve deeper into the roles of local and global structural knowledge, we train THTN excluding local structural (LS) components (specifically, $lse$, $ce$, $lc$, and $kc$) and omitting global structural (GS) elements (namely, $ue$, $hd$, and $cc$) on all dataset. Outcomes in Table 3 illustrate that local structural information significantly influences THTN efficacy more than its global counterpart. Given that hypergraphs inherently harness higher-order global structural details from the graph, the omission of global encodings in THTN does not critically hinder its performance. On the other hand, converting a graph to a hypergraph can result in a loss of local structural details, underscoring the importance of integrating local structural insights into hypergraph learning.
@@
ii. Impact of different design schemes To better understand the impact of each design scheme, we conduct ablation studies on a small and a large dataset, Cora and PubMed, respectively. We consider running THTN without (W/O) considering a specific design scheme each time. The results are presented in Table 4, which implies the impact of the individual design scheme on the model performance. For instance, removing positional encoding ($pe$) results in an accuracy of 85.51% for the Cora dataset and 86.22% for the PubMed dataset. This table shows that for both datasets, the greatest performance degradation occurs when the local structure encoding ($lse$) is excluded from the input node feature. When we create hyperedge from a community, we lose the local connection information between the nodes, which might be very important for the hyperedges. $lse$ preserves local connection information by applying GCN to the input graph. The significant decrease in performance without $lse$ highlights its crucial contribution to the model’s overall performance.
@@
iii. Impact of Feature Normalization and Concatenation As illustrated in Equation 4, our method combines various structural and spatial features (i.e., encodings) with the initial features of the nodes by summing them. To investigate the effect of normalizing these features before combining them with the node’s initial features, we conduct an additional experiment. In this experiment, we normalize the features prior to their integration. The results of this experiment are presented in Table 5 and are labeled as THTNN. Additionally, we perform another experiment to evaluate the impact of concatenating the features instead of summing them. The outcomes of this approach are also shown in Table 5 and are referred to as THTNC. A comparative analysis of the different versions of THTN presented in the table reveals that each version performs better on different datasets. However, the original THTN method generally outperforms both the normalized variant (THTNN) and the concatenated variant (THTNC).
iv. Impact of global nodes We investigate the effect of global nodes ($n_{g}$) on the model’s performance, illustrated in Figure 3. As per Figure 3 (a), the Cora dataset reveals a rising trend in accuracy as the global nodes $n_{g}$ increase, peaking at 88.48% with $n_{g}=3$, followed by a decline. Analogously, Figure 3 (b) presents the Citeseer dataset, where an increment in global nodes prompts an upsurge in accuracy, achieving a maximum of 78.09% at $n_{g}=1$ before the accuracy trend reverses. Similarly, we get the best performance for the PubMed, DBLP, LastFMAsia, and Wiki datasets for $n_{g}=4$, $n_{g}=5$, $n_{g}=4$, and $n_{g}=4$ respectively. Overall, this trend suggests that an optimal number of global nodes can effectively incorporate relevant global information. However, exceeding this optimal number may introduce excessive parameters and reduced generalization ability. Thus, selecting an appropriate number of global nodes is crucial for optimal performance.
@@
Model Cora Citeseer PubMed DBLP LFMA Wiki
THTN 88.48 78.09 89.01 84.38 88.58 79.15
THTNN 86.88 76.23 88.41 85.46 87.26 78.13
THTNC 89.13 77.83 87.18 79.32 87.78 75.25
@@ 5. Conclusion
In this paper, we introduce the Topology-guided Hypergraph Transformer Network (THTN) as a novel hypergraph neural network model for node classification. While existing hypergraph transformer models mainly rely on semantic feature-based attention, they ignore topological information. The THTN has addressed this limitation by presenting a structural and spatial encoding module along with a structure-aware attention mechanism. By considering both semantic and structural information, THTN provides a comprehensive approach to node representation learning. Experimental findings underscore THTN’s superior node classification capabilities relative to benchmarks, highlighting its potential to revolutionize hypergraph analytics.